OpenStudy (anonymous):

If anyone can help me out with this equation, I'd appreciate it. -6/7+z=-8/15

OpenStudy (anonymous):

z=6/7-8/15 z=(6*15 -8*7)/7*15 calculate...

OpenStudy (anonymous):

Help me out here. I have 34/90 there's no simplifying necessary

OpenStudy (anonymous):

Please I need an answer right away.